### v1

# Progressive Growing of Patch Size: Curriculum Learning for Accelerated and Improved Medical Image Segmentation

Abstract
In this work, we introduce Progressive Growing of Patch Size, an automatic curriculum learning approach for 3D medical image segmentation. Our approach progressively increases the patch size during model training, resulting in an improved class balance for smaller patch sizes and accelerated convergence of the training process. We evaluate our curriculum approach in two settings: a resource-efficient mode and a performance mode, both regarding Dice score performance and computational costs across 15 diverse and popular 3D medical image segmentation tasks. The resource-efficient mode matches the Dice score performance of the conventional constant patch size sampling baseline with a notable reduction in training time to only 44%. The performance mode improves upon constant patch size segmentation results, achieving a statistically significant relative mean performance gain of 1.28% in Dice Score. Remarkably, across all 15 tasks, our proposed performance mode manages to surpass the constant patch size baseline in Dice Score performance, while simultaneously reducing training time to only 89%. The benefits are particularly pronounced for highly imbalanced tasks such as lesion segmentation tasks. Rigorous experiments demonstrate that our performance mode not only improves mean segmentation performance but also reduces performance variance, yielding more trustworthy model comparison. Furthermore, our findings reveal that the proposed curriculum sampling is not tied to a specific architecture but represents a broadly applicable strategy that consistently boosts performance across diverse segmentation models, including UNet, UNETR, and SwinUNETR. In summary, we show that this simple yet elegant transformation on input data substantially improves both Dice Score performance and training runtime, while being compatible across diverse segmentation backbones.
keywords:
3D Medical Image Segmentation , Patch-based Segmentation , Patch Sampling , Curriculum Learning , Class Imbalance
journal: Medical Image Analysis
[…]
[aff1]organization=School of Computation, Information and Technology, Technical University Munich, city=Munich,country=Germany
[…]
[aff2]organization=Department of Radiation Oncology, TUM School of Medicine, TUM University Hospital rechts der Isar, Technical University of Munich, city=Munich, country=Germany
[…]
[aff3]organization=Institute of Machine Learning in Biomedical Imaging, Helmholtz Munich,city=Munich, country=Germany
[…]
[aff7]organization=Institute of Radiation Medicine, Helmholtz Center Munich, city=Munich, country=Germany
[…]
[aff4]organization=Munich Center of Machine Learning (MCML),city=Munich, country=Germany
[…]
[aff5]organization=School of Biomedical Engineering and Imaging Sciences, King’s College London,city=London, country=UK
[…]
[aff6]organization=Departament de Matemàtiques i Informàtica, Barcelona Artificial Intelligence in Medicine Lab (BCN-AIM), Universitat de Barcelona,city=Barcelona, country=Spain
{graphicalabstract}[Uncaptioned image]
{highlights}
We propose a novel input-length curriculum based on patch size. By simply adapting the sampling strategy, it can both accelerate training convergence and enhance final segmentation performance over standard constant patch size training in patch-based 3D medical image segmentation. The curriculum can be applied in a performance mode or in a resource-efficient mode.
In performance mode, the curriculum consistently outperforms constant patch size sampling across all tested datasets, achieving statistically significantly higher segmentation performance while also being more resource-efficient.
In resource-efficient mode, the curriculum achieves segmentation performance comparable to constant patch size sampling across all tested datasets, while substantially reducing both runtime and computational cost.
The proposed sampling strategy is particularly effective for highly imbalanced segmentation tasks, where volumes consist mainly of background voxels, as for lesion segmentation, by implicitly improving class balance at training start.
1 Introduction
Most research efforts in the area of medical image segmentation focus on the development of new architectural concepts, including convolution-based [1, 2], transformer-based [3, 4], and hybrid approaches [5, 6]. In contrast, comparatively little attention has been given to the training process itself, where models are still predominantly trained using random data sampling strategies. Inspired by human learning, Bengio et al. [7] introduced the concept of curriculum learning, where models first learn simpler tasks before progressing to more difficult ones. By ordering samples from easy to hard, Bengio et al. [7] demonstrated faster convergence and improved performance relative to random sampling. Such techniques also have the potential to reduce the substantial computational costs of training, both in terms of time and energy consumption, leading to a beneficial environmental impact [8].
Several approaches have been proposed to design curricula for medical imaging. Some rely on human annotations or expert knowledge to define task-specific measures of difficulty [7, 9, 10]. For example, task difficulty might be linked to class membership, as in fracture classification [9], or defined using inter-rater agreement [9, 10]. However, expert annotations considerably increase the costs, making those approaches often infeasible. More general strategies estimate task difficulty automatically or synthetically change data complexity during training. Sample loss has been used as a proxy for difficulty, enabling hard-negative mining and adaptive oversampling, as shown in lung cancer segmentation [11]. In MRI-based brain tumor segmentation, Havaei et al. [12] introduced a curriculum that improves robustness to missing modalities by randomly dropping channels with increasing probability during training. Karras et al. [13] proposed to progressively grow a generative adversarial network layer by layer, effectively increasing task difficulty as output resolution doubled. This approach was used for natural image generation and showed improved convergence, reduced training time, and an overall better model performance. Zhao et al. [14] extended this idea to semantic segmentation of cervical nuclei by progressively growing the UNet architecture from its bottleneck.
Another way to define task difficulty is by sample length. In natural language processing, curricula based on sentence length have been used to accelerate the training of large language models such as BERT and GPT-2 [15, 16, 17, 18] and have been adopted into the high-performance training library DeepSpeed [19]. In computer vision, an analogous measure is image size. Several works employ curricula that (progressively) increase input resolution [20, 21, 22, 18] during training. These methods consistently improve convergence, enabling either more resource-efficient training with comparable performance [20, 21] or higher final accuracy within the same training budget [22]. Several curricula were applied for visual backbone pretraining tasks [22, 21], by pretraining on ImageNet classification [21, 20] or contrastive pretraining [22, 23, 24]. In contrast, there is limited work in applying the sample-length curriculum directly to segmentation tasks [25], thus semantic segmentation mostly benefits from curriculum learning only indirectly via transferring pretrained weights [22, 21].
In this work, we propose a Progressive Growing of Patch Size curriculum, which is a novel sample-length curriculum for 3D medical image segmentation, defined through patch size. Our approach starts training on small patches, ensuring strong class balance in early training, and progressively increases to large patches, which provide a broader global context. Furthermore, the reduced memory footprint of smaller patches early in training allows for larger batch sizes under the same GPU budget. We argue that our proposed curriculum, built on patch size, is better suited for dense prediction tasks such as semantic segmentation than the Progressive Resolution curriculum applied in multiple methods in the CV domain [20, 21, 22]. Furthermore, our approach is an advancement over the regularly applied constant patch size sampling.
Building on our previously proposed Progressive Growing of Patch Size curriculum at MICCAI 2024 [26], we extended the methodology and analysis in the following way: (i) We extended the originally proposed curriculum by a performance mode, using maximally large batch sizes, enabling full GPU utilization; (ii) We validated the generalization to different backbones by applying it to one fully convolutional network, UNet, and two transformer-based networks, UNETR and SwinUNETR; (iii) We analyzed the stochastic differences in performance in repeated training runs due to stochastic training differences; (iv) We compared our proposed curriculum based on patch size to a similar existing curriculum based on image resolution. The extended methodology and thorough evaluation show that our proposed curriculum, built on patch size, improves segmentation performance in the form of Dice score and at the same time reduces the computational costs of training.
2 Background on 3D Medical Image Segmentation
Three-dimensional (3D) medical image segmentation is fundamentally constrained by the large GPU memory requirements of volumetric data processing. Common strategies to address this limitation include: (1) using low-resolution models [2], (2) applying high-resolution patch-based methods [6, 5, 27, 2, 28], and (3) employing network cascades [29] or multi-scale approaches [30]. Among these, it has been shown that high-resolution patch-based methods generally yield superior performance [29].
Patch sampling strategies play a central role in patch-based methods. nnU-Net adopts a “forced oversampling” technique, where one foreground (FG) patch is sampled from a randomly chosen patient, with FG classes selected with equal probability, while another random patch is sampled from a different patient [2]. In contrast, MONAI implements “Probabilistic Oversampling” which applies probabilistic class sampling [31].
To maximize global context, standard practice is to use the largest possible patch size that fits into GPU memory, typically with a batch size of two [2]. Even with additional computational resources, nnU-Net prioritizes maximizing patch size over increasing batch size, as exemplified in its larger encoder variants such as nnU-Net ResEnc M/L/XL [32].
Figure 1: Overview of the proposed Progressive Growing of Patch Size curriculum, illustrated for lung cancer segmentation (cancer regions highlighted with yellow bounding boxes). Training begins with the minimal patch size ($H_{min},W_{min}$) and progressively increases the patch dimensions stepwise until the final maximal patch size ($H_{max},W_{max}$) is reached. Smaller patch sizes provide a higher ratio of foreground to background voxels, which decreases as the patch size grows. During inference, the maximum patch size is used to capture maximal global context. Figure is adapted from [26].
3 Methods
This section describes the proposed methods and is split into three parts. First, we discuss the theoretical design of the proposed curriculum learning. After that, we introduce the efficient curriculum mode and the performance mode. Furthermore, we outline the nnU-Net framework implementation of the curricula.
3.1 Progressive Growing of Patch Size
The central principle of our proposed Progressive Growing of Patch Size (PGPS) is to begin the training with the smallest processable patch size and progressively increase to larger patches. The rationale for starting with smaller patches lies in task difficulty: smaller patches inherently yield a more balanced foreground-to-background ratio, whereas this balance diminishes as patch size increases. In the theoretical case of a single-voxel patch, a batch containing one foreground and one background patch would achieve perfect balance, with half of the voxels belonging to the foreground and half to the background. As the patch size grows, however, the distribution converges toward the full volume statistics, which are typically dominated by background voxels.
Patch size increments are applied in the smallest feasible steps, with each axis adjusted independently. By gradually increasing the patch size in minimal increments, transitions between training stages are smoothed, as the segmentation tasks at successive patch sizes remain closely related. This design provides the network with a structured sequence of progressively more challenging tasks. An overview of this curriculum is illustrated in Figure 1.
Both the minimal patch size and the amount of increment depend on the underlying network architecture. The progression continues until the largest possible patch size is reached, which is typically constrained by the GPU memory capacity or the network’s default patch size. During inference, the largest patch size is employed to maximize global context, which usually achieves the highest segmentation performance [2].
3.2 Curriculum Modes: Efficiency vs. Performance
We propose two modes of the Progressive Growing of Patch Size (PGPS) curriculum, tailored to different objectives. PGPS-Efficiency minimizes training runtime while maintaining performance comparable to standard constant patch size sampling. PGPS-Performance aims to maximize segmentation performance. Intermediate configurations between these two extremes are also possible. Figure 2 illustrates both modes for a multi-organ segmentation task.
PGPS-Efficiency: In PGPS-Efficiency, the batch size is kept constant throughout training. Since the patch size is smaller at early stages, the number of computations is reduced, leading to a substantial reduction in training time. This is illustrated in Figure 2 in green.
PGPS-Performance: In PGPS-Performance, the GPU memory budget is fully utilized by dynamically increasing the batch size when smaller patches are used. This design leverages available resources more effectively, with the goal of achieving the highest possible segmentation performance. This is illustrated in Figure 2 in yellow.
3.3 nnU-Net Framework Implementation
In the following section, we describe the integration details of default constant patch size sampling and the two different proposed PGPS sampling modes into the nnU-Net framework [32], which is the state-of-the-art 3D medical image segmentation framework.
3.3.1 Baseline: Constant Patch Size Sampling
The baseline sampling method utilized in this study is fixed or constant patch size (CPS) sampling. This sampling strategy corresponds to the conventional configuration of training a (patch-based) segmentation network.
We have chosen nnU-Net [32] as the underlying framework, as it provides state-of-the-art performance and is auto-configuring to downstream tasks. The default nnU-Net configuration is designed to achieve maximum global context by employing the largest possible patch size that fits within nnU-Net’s typical GPU budget of 8GB.
Given a patch-based setting, where the maximal processable patch size is smaller than the target volume size, nnU-Net will by default apply a batch size of two. In this setting, two patients are randomly sampled, while from the first patient, a foreground (FG) patch will be created, and from the second patient, a random patch will be sampled. This default 50% FG patch ratio for a batch is designed to improve robustness in the presence of class imbalance [2]. If multiple FG classes are present, one class will be picked randomly with equal probabilities. If nnU-Net sets a batch size larger than two, the framework employs a foreground patch ratio of 33%.
3.3.2 PGPS Curricula
We integrate the proposed Progressive Growing of Patch Size curricula within the nnU-Net framework. This integration involves maintaining the default nnU-Net settings while only adjusting the patch size and the batch size for PGPS-Performance.
For the PGPS curricula, we have to define the minimal processable patch size and, given the maximal patch size, find all processable intermediate patch sizes that the network can process. In the case of nnU-Net, which applies a fully convolutional UNet backbone, the smallest processable patch size is dependent on the number of downsampling operations. The minimum patch size length depends on the number of downsampling operations for a UNet, and is given by $2^{num\_pool}$, where num_poolnum\_pool is the number of pooling operations for each axis. All intermediate patch sizes have to be fully divisible by $2^{num\_pool}$. The maximal patch size is the patch size nnU-Net would by default apply to the given task.
Figure 2: Input tensors for the two proposed modes of the Progressive Growing of Patch Size (PGPS) curriculum, illustrated for a multi-organ segmentation task with example foreground patches (top) and corresponding label maps (bottom). In both modes, the patch size is progressively increased during training. In PGPS-Efficiency (green), the batch size remains constant, resulting on average in smaller input tensors. In PGPS-Performance (orange), the available GPU budget is fully utilized by increasing the batch size.
The patch size is increased in a step-wise linear fashion from the minimal possible patch size that the UNet allows until the patch size reaches the nnU-Net’s default patch size. To have a smaller input tensor size growth, we only increase the patch size for one axis per step. We always select the axis with the lowest numerical value to increase the patch size. This scheme results in a higher average batch size during training, compared to naively sequentially increasing each axis. We train each patch size stage with the same number of epochs.
PGPS-Efficiency: For PGPS-Efficiency, only the patch size is adjusted during training, while the batch size is kept constant, using the default nnU-Net batch size. This results in reduced GPU memory utilization for smaller patch sizes.
PGPS-Performance: In PGPS-Performance, both patch size and batch size are adjusted to fully utilize the available GPU memory budget. By default, nnU-Net enforces a foreground patch ratio of 33% for batch sizes larger than two, but switches to 50% when the batch size is two. This setting would introduce a discontinuity in the class balance trajectory. To ensure a smooth progression, we instead fix the foreground patch ratio to 50%, regardless of batch size. As there are different options in how to increase the batch size, we evaluated multiple batch construction strategies regarding efficiency and segmentation performance in A. We identified that creating all patches from two patients per batch, using one patient for foreground patches and the other for random patches, performs best in terms of efficiency and segmentation performance. This allows the creation of overlapping patches within the same patient. We follow this batching strategy throughout the manuscript.
4 Experiments
In Section 4.1, we introduce a set of widely used public medical image segmentation datasets used for the experiments. In Section 4.2, we evaluate the PGPS curriculum modes, analyze segmentation performance, computational costs (runtime and FLOPs), and convergence properties, and compare them to the constant patch size baseline. We present a correlation analysis between performance differences and task-specific characteristics in Section 4.3. A comparison between Progressive Resolution [20] and our proposed method is provided in Section 4.4. In Section 4.5, the generalization of PGPS to different vision backbones is investigated. Finally, the stochastic variability of the model performance across repeated trainings with the three sampling strategies is examined in Section 4.6.
4.1 Dataset Descriptions
To test the curriculum methods, a wide range of publicly available segmentation datasets was utilized. In total, 15 different 3D medical semantic segmentation datasets were gathered. The dataset mix included the Medical Segmentation Decathlon [33], comprising 10 distinct datasets covering lesion, organ, and vessel tasks. Additionally, the ToothFairy2 dataset [34], TotalSegmentatorV2 [35], KiTS23 [36], AMOS22 CT/MRI task [37] and BTCV [38] were incorporated. These datasets encompassed different segmentation tasks, including lesions, various organs, bones, muscles, teeth, implants, and nerves. Furthermore, the datasets represented different 3D modalities such as CT, Conebeam-CT, and MRI, and varied in size from small to large, ranging from 20 to 1200 samples. They presented a diverse range of class numbers, from 2 (binary) to 117 classes. TotalSegmentatorV2 was used as a single 117-class task by combining all five subtasks, namely organs, vertebrae, cardiac, muscles, and ribs.
4.2 Benchmarking of Methods on 15 Datasets
We evaluated segmentation performance, training runtime, and computational cost (FLOPs) of the PGPS curricula relative to default constant patch size (CPS) training. The default nnU-Net preprocessing and training pipeline was applied, using the 3D high-resolution patch-based variant for all tasks. For ToothFairy2 and TotalSegmentatorV2, mirroring data augmentation was disabled, as prior studies demonstrated improved results without mirroring [39, 35]. For both datasets, we also evaluated the segmentation performance for nnU-Net training without mirroring data augmentation in B.
4.2.1 Segmentation Performance
To benchmark CPS, PGPS-Efficiency, and PGPS-Performance, we followed the standard nnU-Net evaluation procedure: models were trained using 5-fold cross-validation, and Dice Scores were averaged across all foreground classes.
To assess whether segmentation performance differs significantly between CPS and the PGPS curricula across the 15 datasets, paired Wilcoxon signed-rank tests were performed on the average foreground Dice Scores. For this, we built one sample per dataset by pairing curriculum and CPS score.
4.2.2 Training Runtime Tracking
Training time was monitored, excluding validation. Since experiments were conducted on a cluster with heterogeneous GPU and CPU configurations, we propose a virtual relative runtime to enable fair comparisons. For the PGPS curricula, time per epoch was recorded. We then compute the average runtime for all maximal patch size epochs, which then represents the runtime of a single CPS epoch. Given that, we can then compute the virtual runtime of a full CPS training under the same hardware settings. We report the median over all 5-folds for the runtime.
4.2.3 FLOPs Tracking
The number of Floating Point Operations (FLOPs) is directly proportional to the size of input tensors. We recorded the total number of iterated voxels during training for each strategy and normalized them to the CPS value, yielding the relative difference in computational cost between PGPS curricula and CPS.
4.2.4 Convergence Analysis
We analyzed the convergence behavior of the three sampling strategies. For each strategy, models were trained across all 15 datasets with varying training lengths: 1%, 10%, 25%, 50%, and 100% of total training iterations. This is done by reducing the number of iterations per epoch relative to the default 250 training iterations for nnU-Net. Evaluation at each training length was performed using 5-fold cross-validation.
4.3 Task Characteristics Analysis
To identify dataset characteristics associated with segmentation performance improvements, we correlated relative performance gains with the following task measures: (1) number of semantic classes, (2) training dataset size, (3) patch-to-volume coverage (the proportion of a full volume seen in a single patch), and (4) class imbalance, measured by the frequency of the smallest class.
Spearman correlation tests were conducted for each task characteristic against the relative Dice Score improvement between CPS and PGPS curricula. Separate tests were performed for each training length (1%, 10%, 25%, 50%, and 100%).
Additionally, we recorded the foreground ratio of input tensors, the number of unique classes seen per iteration, and the patch size ratio across all three sampling strategies.
4.4 Progressive Resolution Curriculum
Another input-length curriculum in computer vision is progressive resizing, also called progressive resolution [22, 20], which gradually increases input resolution from low to high. This approach also modifies input length, transitioning from small to large input tensors. To compare our PGPS curricula with progressive resolution, we evaluated both on BTCV, AMOS22, KiTS23, and MSD Lung Cancer, covering two highly imbalanced lesion tasks and two multi-organ tasks.
We implemented the Progressive Resolution curriculum within the nnU-Net framework. Input tensor sizes were matched to those used for the PGPS curricula; however, instead of adjusting crop sizes, the default nnU-Net patch size was resampled to fit PGPS’ target tensor size for each phase. Increasing the batch size, as done in PGPS-Performance, is currently not efficient within nnU-Net for progressive resolution; therefore, we only evaluated Progressive Resolution using a fixed batch size. Segmentation performance of both curricula was compared using Dice Scores in a 5-fold cross-validation.
[…]
Dataset Dice Score [%] $\uparrow$ $\blacktriangle x\%$ Rel. Dice Score $\uparrow$ Rel. Runtime [%] $\downarrow$ Rel. FLOPs [%] $\downarrow$
CPS PGPS-Eff PGPS-Perf PGPS-Eff PGPS-Perf PGPS-Eff PGPS-Perf PGPS-Eff PGPS-Perf
MSD Brain 74.12 74.29 74.15 $\blacktriangle$ 0.23 $\blacktriangle$ 0.04 42 83 34 84
[…]
MSD Liver 78.74 78.76 80.60 $\blacktriangle$ 0.03 $\blacktriangle$ 2.36 51 85 38 84
MSD Hippocampus 88.96 89.12 89.15 $\blacktriangle$ 0.20 $\blacktriangle$ 0.22 67 112 33 92
MSD Prostate 73.13 75.26 76.31 $\blacktriangle$ 2.91 $\blacktriangle$ 4.35 43 87 30 87
MSD Lung 70.00 72.30 72.77 $\blacktriangle$ 3.29 $\blacktriangle$ 3.96 41 82 31 89
[…]
MSD Colon 48.41 50.83 51.02 $\blacktriangle$ 5.00 $\blacktriangle$ 5.39 38 87 28 90
[…]
KiTS23 86.02 87.22 86.46 $\blacktriangle$ 1.40 $\blacktriangle$ 0.51 45 86 27 84
[…]
ToothFairy2 76.92 77.15 77.00 $\blacktriangle$ 0.31 $\blacktriangle$ 0.11 34 85 26 88
[…]
Norm. Avg. 100.00 100.47 101.26 $\blacktriangle$ 0.47 $\blacktriangle$ 1.26 $44\pm 9.5$ $89\pm 8.7$ $33\pm 2.6$ $88\pm 2.4$
4.5 Different Architectures
The proposed curriculum can be applied to any vision backbone that supports flexible input sizes. To assess its applicability beyond CNNs, we further evaluated the curriculum on transformer-based hybrid architectures, specifically UNETR [5] and SwinUNETR [6], both combining a transformer encoder with a CNN decoder.
In transformer-based models, architectural constraints define the minimal feasible patch size and increment: for UNETR, this is determined by the internal patchify size (distinct from the patch size used in patch-based training), while for SwinUNETR it is defined by the Swin window size. As transformers typically rely on fixed input dimensions due to positional embeddings, we interpolated the positional embeddings to match the varying patch sizes as in [3].
Both architectures were integrated into the nnU-Net framework, trained following the standard nnU-Net training, and evaluated on the BTCV dataset using a 5-fold cross-validation. We did train all backbones from scratch.
4.6 Stochastic Variability in Training
Neural network training is inherently stochastic, a property that is amplified in patch-based pipelines, where both the sampled volume and patch location are randomly selected (within oversampling constraints). Consequently, repeated training runs can yield different models and segmentation performance.
To quantify this variability, we repeated training on a single fold of the 5-fold cross-validation five times for each strategy. Experiments were conducted on the BTCV multi-organ segmentation dataset and the highly class-imbalanced MSD Lung Tumor dataset. Training lengths of 1%, 10%, and 100% of the default iteration count were analyzed.
To assess the stochastic impact of sampling, we build triplets of each strategy’s segmentation performance outcome and evaluate all 125 possible outcome combinations per training length. As each single outcome could be a potential result, we are interested in which sampling strategy ranks best in each possible comparison. Thus, we count the number of scenarios in which each strategy outperformed both other strategies.
5 Results
5.1 Benchmarking of Methods on 15 Datasets
We evaluated the Progressive Growing of Patch Size (PGPS) curricula relative to the conventional constant patch size (CPS) baseline across 15 popular 3D medical image segmentation tasks. Key measures included segmentation performance, training runtime, and computational cost in terms of FLOPs. Additionally, we also analyze the training convergence.
5.1.1 Segmentation Performance
Table 1 summarizes the performance benchmarking of the baseline compared to the curricula approaches on 15 datasets. Across the 15 datasets, PGPS-Performance achieved the highest average Dice Score, yielding a relative improvement of 1.26% over CPS. It outperformed CPS in every task, while ranking as the best-performing strategy in 12 of the 15 tasks. In all 15 datasets, PGPS-Performance was among the top two strategies, and only in the three tasks, KiTS23, ToothFairy2, and MSD Brain Tumor, PGPS-Performance was surpassed by PGPS-Efficiency.
PGPS-Efficiency also delivered modest overall performance gains, with a relative improvement of 0.47% over all 15 tasks compared to CPS. It outperformed CPS in 8 tasks and was leading in 3 tasks. However, in 7 tasks, it had the lowest Dice Score.
A two-sided paired Wilcoxon signed-rank test comparing PGPS-Performance to CPS across all 15 datasets revealed a significant improvement for PGPS-Performance ($p\approx 0.0001$). Additionally, PGPS-Performance significantly outperformed PGPS-Efficiency ($p\approx 0.0103$), confirming it as the superior strategy in terms of segmentation performance.
A two-sided paired Wilcoxon signed-rank test between PGPS-Efficiency and CPS showed no significant difference ($p\approx 0.6788$), indicating comparable segmentation performance between these two strategies.
5.1.2 Training Runtime Tracking
Table 1 shows the relative runtime for all 15 datasets. PGPS-Performance required approximately 89% of CPS’s runtime, ranging from 82% to 112% across tasks. Tasks for which PGPS-Performance yielded longer training times than CPS, i.e. MSD Hippocampus and TotalSegmentatorV2 had the two largest start batch sizes of 1575 and 784, respectively. In total, the 15 tasks had an average batch size of around 399, ranging from 128 to 1575. PGPS-Efficiency substantially reduced training time, requiring only about 44% of CPS runtime (range: 34%–67%), achieving the fastest training while maintaining comparable segmentation performance.
Figure 3: Segmentation performance convergence of CPS, PGPS-Performance, and PGPS-Efficiency across training iterations. Dice Scores are tracked across 15 segmentation tasks for different training lengths (1%, 10%, 25%, 50%, and 100% of nnU-Net’s default total training iterations). On average, PGPS-Performance exhibits the fastest convergence, while PGPS-Efficiency converges more slowly, due to the smaller input tensors that result in superior training speed. The final performance is, on average, best for PGPS-Performance, followed by PGPS-Efficiency and CPS. The average convergence over all 15 datasets (right) has been computed by normalizing each task by its maximum performance and then averaging across all tasks.
5.1.3 FLOPs Tracking
As shown in Table 1, PGPS-Efficiency utilized only about 33% of the FLOPs of CPS on average, with a range of 26%–38%. Although PGPS-Performance is designed to maximize the used GPU memory, it required slightly fewer FLOPs than CPS, averaging 88% (range: 84%–92%). The reduced computational cost arises from the discrete nature of batch and patch sizes, which prevents full usage of the GPU budget in every epoch, resulting in lower FLOPs.
5.1.4 Convergence Analysis
Figure 3 shows the convergence of segmentation performance across all 15 tasks over the number of iterations.
At 1% training length, PGPS-Performance already achieved the highest Dice Score in 13 tasks, ranking second to CPS only in MSD Spleen and KiTS23. At 10% and 25%, PGPS-Performance outperformed both other strategies across every task. At 50%, PGPS-Performance outperformed both other strategies except for MSD Heart. However, the minimal differences in the results across methods and between training at 50% and 100% suggest that the performance is saturated in the MSD Heart task.
Overall, PGPS-Performance consistently converged faster than both CPS and PGPS-Efficiency, achieving superior results in the majority of training-length scenarios for each task.
In contrast, PGPS-Efficiency exhibited slower convergence than both other sampling strategies, as shown in Figure 3, but required substantially fewer computations, as seen in Table 1, processing only about one-third of the voxels. Nevertheless, over the course of full training, PGPS-Efficiency reached comparable, and in some cases even superior segmentation performance compared to CPS.
5.2 Task Characteristics Analysis
To better understand the improved convergence for PGPS curricula, we tracked several task characteristics during model training for the BTCV dataset. Tracked characteristics included foreground class balance, number of unique classes per iteration, and patch size ratio. These are illustrated in Figure 4. We find that both PGPS curricula exhibit higher class balance during early training phases compared to CPS, which maintains at a stationary foreground-to-background ratio. The reason is that fewer surrounding background voxels are contained in smaller foreground patches. At the final stage, PGPS curricula converge to the same class balance as CPS due to identical patch sizes. The standard deviation of class balance is larger for PGPS-Efficiency than for PGPS-Performance. PGPS-Performance shows a discontinuous drop in class balance during later training stages, attributed to the discrete nature of batch and patch sizes.
Figure 4: Training characteristics of PGPS curricula for the BTCV dataset with 14 classes. Left: PGPS curricula improve foreground-background voxel balance compared to CPS due to smaller patch sizes. Center: PGPS-Performance achieves the highest average number of unique classes per batch/iteration due to larger batch sizes, while PGPS-Efficiency has the lowest. Right: CPS maintains a constant input tensor size; PGPS-Efficiency increases tensor size exponentially, while PGPS-Performance exhibits occasional size drops due to the discrete nature of patch and batch sizes.
Tracking the number of unique classes per iteration for the 14-class BTCV task, PGPS-Performance produces batches containing nearly all semantic classes ($\sim 100\%$) during early patch size stages, converging to the stationary CPS value ($\sim 82\%$) in later stages. Small drops in unique classes per iteration are again observed due to the discrete batch and patch sizes. PGPS-Efficiency begins with the smallest fraction of unique classes per iteration ($\sim 18\%$), which gradually increases across patch size stages, eventually converging to the CPS value. The input tensor sizes of PGPS-Performance are the same as for CPS, but have occasional drops due to the discrete nature of batch and patch size. These drops result in the shorter training runtime observed for PGPS-Performance relative to CPS. PGPS-Efficiency exhibits a monotonic, stepwise exponential increase in patch size ratio.
PGPS-Performance: Spearman tests did not yield any significant correlation between PGPS-Performance relative improvements and the dataset characteristics of the number of semantic classes, dataset size, or patch-to-volume coverage. The only significant correlation was negative, occurring for class imbalance measured via the smallest class frequency at training lengths of 1% ($\rho\approx-0.7429$, $p\approx 0.0015$) and 10% ($\rho\approx-0.5464$, $p\approx 0.0351$). This indicates that PGPS-Performance is particularly beneficial for highly imbalanced tasks.
PGPS-Efficiency: A significant positive correlation was observed between class imbalance (smallest class frequency) and relative performance at training lengths of 1% ($\rho\approx 0.5321$, $p\approx 0.0412$), 10% ($\rho\approx 0.7679$, $p\approx 0.0008$), and 25% ($\rho\approx 0.6071$, $p\approx 0.0134$). Highly imbalanced tasks show reduced performance in short training scenarios, but PGPS-Efficiency eventually outperforms CPS at full training. This trend is evident for lesion tasks such as MSD Brain, MSD Liver Tumor, MSD Prostate, MSD Lung Tumor, MSD Colon, and KiTS23. Additionally, patch-to-volume coverage correlates significantly with relative performance at training lengths of 10% ($\rho\approx 0.5157$, $p\approx 0.0491$) and 25% ($\rho\approx 0.5318$, $p\approx 0.0382$).
Table 2: Dice Scores of Progressive Resolution (Prog. Res.) and Progressive Growing of Patch Size (PGPS) curricula. PGPS increases patch size during training, whereas Progressive Resolution increases patch resolution. PGPS-Efficiency consistently outperforms Progressive Resolution, with the largest gains observed in highly imbalanced lesion tasks. Across all datasets, Progressive Resolution was the lowest-performing strategy. Results are reported as mean Dice Score (%) of a 5-fold Cross-Validation. Bold: best-performing strategy.
Dataset Prog. Res. PGPS-Eff
MSD Lung Tumor 69.42 72.30
KiTS23 84.63 87.22
BTCV 82.93 83.05
AMOS22 87.98 88.10
5.3 Progressive Resolution Curriculum
We evaluated the Progressive Resolution curriculum, an input-length strategy widely applied in computer vision [20, 21, 22], by adapting it to 3D patch-based medical image segmentation. Table 2 shows Dice Scores for Progressive Resolution and PGPS-Efficiency. We found that our proposed PGPS-Efficiency consistently outperformed Progressive Resolution.
In the lesion segmentation task MSD Lung Tumor, Progressive Resolution lagged behind PGPS-Efficiency by 2.9 Dice points (69.42% vs. 72.30%), and in KiTS23 by 1.2 points (84.63% vs. 87.22%). These results highlight that Progressive Resolution struggles in highly imbalanced settings, where PGPS-Efficiency provides a clear advantage.
For multi-organ segmentation tasks, such as BTCV and AMOS22, the performance gap was smaller. Progressive Resolution achieved 82.93% Dice on BTCV, close to PGPS-Efficiency (83.05%). Similarly, in AMOS22, it reached 87.98%, compared to 88.10% for PGPS-Efficiency. Although differences in these multi-organ tasks were less pronounced, Progressive Resolution was still the lowest-performing strategy in every case.
Figure 5: Segmentation performance for repeated training of CPS, PGPS-Efficiency, and PGPS-Performance across different training lengths (1%, 10%, and 100% of nnU-Net’s total training iterations). Each experiment was repeated five times on the same data split with different random seeds. Variability decreases with longer training, while mean performance increases. The highly class-imbalanced MSD Lung Tumor dataset exhibits higher variance than the multi-class BTCV dataset. PGPS-Performance does improve convergence and reduces performance variability between repeated runs. PGPS-Efficiency resulted in slower convergence than CPS and PGPS-Performance and had the highest performance variability. PGPS-Performance did win over both other strategies in 100% of the 125 possible combinations for MSD Lung Tumor, while for BTCV, PGPS-Performance won over both other strategies in 58.4% of all combinations and outperformed CPS and 60% of combinations.
5.4 Different Architectures
Table 3: Impact of PGPS-Efficiency (PGPS-Eff) and PGPS-Performance (PGPS-Perf) on different backbones for the BTCV task. Results are reported as mean Dice Score (%) across five folds. PGPS-Performance consistently improved performance over CPS across all architectures. PGPS-Efficiency converged for UNet and SwinUNETR but failed for UNETR due to gradient instability. Results are reported as mean Dice Score (%) of a 5-fold Cross-Validation. Bold: best-performing strategy. Underlined: second-best-performing strategy.
Backbone CPS PGPS-Eff PGPS-Perf
UNet (nnU-Net) [2] 83.37 83.05 83.81
UNETR [5] 71.20 0.00 72.76
SwinUNETR [6] 71.62 71.98 75.39
We further assessed the generalizability of PGPS by applying it to transformer-based backbones (UNETR, SwinUNETR). Results for the BTCV dataset are reported in Table 3 and compared to nnU-Net’s default UNet backbone.
PGPS-Performance consistently improved segmentation performance across all architectures. Compared to CPS, Dice Scores increased by +0.44+0.44 points for nnU-Net’s fully convolutional UNet (83.37% $\rightarrow$ 83.81%), +1.56+1.56 points for UNETR (71.20% $\rightarrow$ 72.76%), and +3.77+3.77 points for SwinUNETR (71.62% $\rightarrow$ 75.39%). These results indicate that PGPS-Performance provides benefits not only for convolutional but also for transformer-based models.
In contrast, PGPS-Efficiency yielded mixed results. For nnU-Net, performance slightly decreased by 0.32-0.32 Dice points (83.37% $\rightarrow$ 83.05%), and for SwinUNETR it improved marginally by +0.36+0.36 points (71.62% $\rightarrow$ 71.98%). For UNETR, however, all training runs with PGPS-Efficiency diverged due to gradient explosion.
5.5 Stochastic Variability in Training
Figure 5 summarizes the variability in segmentation performance under repeated training. MSD Lung Tumor did show overall higher segmentation performance variability than for the BTCV dataset. Two consistent trends were observed: (i) outcome deviation per strategy decreases as training length increases, and (ii) sampling strategy PGPS-Performance dominates across both datasets and training length scenarios in segmentation performance over CPS and PGPS-Efficiency.
BTCV: PGPS-Performance dominated early training (1–10%), achieving the highest segmentation performance in all comparisons against CPS and PGPS-Efficiency. At full training, it still achieved the majority of wins (58.4%) over CPS and PGPS-Efficiency and the highest Dice Score ($79.57\%\pm 0.37$), followed by CPS ($79.16\%\pm 0.47$) and PGPS-Efficiency ($78.85\%\pm 0.41$).
MSD Lung Tumor: Across all training lengths, PGPS-Performance won every sampling strategy comparison and achieved the highest Dice Score ($78.98\%\pm 1.10$), outperforming CPS ($74.11\%\pm 1.23$) and PGPS-Efficiency ($73.34\%\pm 2.86$). Variability in segmentation performance per sampling strategy was larger at short training lengths but stabilized as training progressed.
6 Discussion
In this work, we introduced a novel curriculum learning strategy that progressively increases patch size during training. The curriculum was implemented in two modes: one optimized for segmentation performance and the other for computational efficiency. We evaluated Dice score performance, training convergence, and computational cost across 15 diverse datasets covering lesion, multi-organ, vessel, muscle, and skeletal segmentation, and compared it to standard constant patch size (CPS) sampling. The performance mode achieved significant Dice score improvements over CPS while simultaneously reducing computational costs in terms of FLOPs and training time. The efficiency mode matches the segmentation performance of CPS, while drastically cutting both FLOPs and training time by more than half. In addition, we successfully applied the curriculum on multiple network backbones, including UNet, UNETR, and SwinUNETR, thereby confirming its broad applicability. In the following, we discuss the key factors driving these performance gains, as well as the limitations and implications of the Progressive Growing of Patch Size (PGPS) approach.
6.1 Factors Contributing to Segmentation Performance Gain
The segmentation performance gains achieved by PGPS curricula can be explained by several factors. First, PGPS improves class balancing, which significantly correlates with Dice Score improvements. The reason for a higher foreground voxel frequency in PGPS is that fewer surrounding background voxels are contained in smaller foreground patches. Furthermore, in PGPS-Performance, the model is exposed to a larger number of unique classes per iteration, although the correlation between class count and performance was not statistically significant. Additionally, the larger batch sizes in PGPS-Performance increase exposure to data augmentations and promote sampling from more diverse patch locations, leading to broader coverage of the data distribution without losing spatial resolution or altering the forced foreground-to-background ratio. This is supported by the near-significant correlation between patch-to-volume coverage and performance improvement ($\rho\approx-0.4978$, $p\approx 0.0590$) at 50% training length for PGPS-Performance. These results are consistent with existing studies on LLM training, indicating that shorter input sequences reduce gradient variance, improving training stability [16].
Interestingly, most correlations were strongest at shorter training lengths. We attribute this to performance saturation at full training, which dampens differences between sampling strategies.
Furthermore, we hypothesize that the aforementioned factors have contributed to the overall performance gain of PGPS over CPS training, but disentangling those factors will be subject to future research. While PGPS-Efficiency cut training runtime strongly, while keeping comparable segmentation performance to CPS, PGPS-Performance was able to significantly outperform standard CPS training on average on 15 datasets, improving segmentation performance on each dataset, while also slightly reducing computational costs.
6.2 PGPS on Different Backbones
PGPS-Performance improved the segmentation performance for all different tested backbones. Therefore, we hypothesize that limiting the contextual information available to the network, by training on smaller image patches, facilitates more stable and efficient learning, similar to findings reported for large language models in [16]. When trained from scratch, Transformer-based architectures generally lag behind CNNs due to their lack of strong inductive biases [40]. While large-scale pretraining is a common solution to mitigate this limitation [3], several approaches have been proposed to introduce inductive bias directly into Transformers, such as CNN-based teacher distillation [40] or architectural modifications like shifted window attention [41, 4].
Our results demonstrate that the proposed PGPS-Performance curriculum consistently improves segmentation performance across all evaluated architectures, including UNet, UNETR, and SwinUNETR. In contrast, the PGPS-Efficiency variant converged successfully for UNet and SwinUNETR but failed for UNETR. We hypothesize that the stronger inductive biases in UNet and SwinUNETR enable effective learning even under reduced input context, whereas UNETR appears more sensitive to limited information, leading to training instabilities. PGPS-Performance seems to enhance the training signal compared to constant patch size training, thereby improving Dice scores for UNETR as well. Another potential factor for UNETR’s instability under PGPS-Efficiency is that simple interpolation of positional embeddings may be suboptimal, while the other two architectures do not rely on such embeddings.
In summary, we hypothesize that training with PGPS-Performance facilitates the learning process by improving the quality of the training signal, particularly for architectures lacking strong inductive bias, such as Transformers. Although we did not explore pretrained Transformer weights in this work, an open research question remains how transfer learning and pretrained representations interact with PGPS-based curricula.
6.3 Comparison to Progressive Resolution
Our results demonstrate that PGPS consistently outperforms Progressive Resolution in segmentation tasks. This is likely due to improved class balancing with PGPS, while Progressive Resolution does not affect the class balance and thereby matches the class balance of CPS. The strong performance drop of Progressive Resolution in highly imbalanced tasks further supports that claim. Progressive Resolution is thus only suited for global image-level prediction tasks such as classification [20], image generation [13], contrastive learning [24], and CLIP alignment [23]. By contrast, PGPS is more effective for dense prediction tasks like segmentation and potentially object detection, particularly in class-imbalanced scenarios.
6.4 Statistical Variance and Model Validation
A further challenge is the variance in reported performances due to statistical variance in deep learning training. For example, Dice scores on the BTCV task vary widely, even under the same data splits: 83.37% (ours), 83.56% (MedNeXt [27]), and 83.08% (nnU-Net Revisited [32]). Surprisingly, even the stronger and larger nnU-Net ResEnc variants (M: 83.31%, L: 83.35%, XL: 83.28%) reported in [32] are outperformed by smaller default nnU-Net models in MedNeXt [27] and in our study. While the extent to which sampling alone explains these discrepancies remains unclear, differences in preprocessing (e.g., different sample origin cache in nnU-Net) may also contribute.
Furthermore, to assess the statistical variance of model training, we report nnU-Net validation performance from related work that were trained on one of our benchmarking datasets, and compare them to the proposed PGPS framework in  C. We found that our proposed curriculum resulted in better-performing models than those reported in the literature, which were trained via conventional CPS sampling.
In [32], Isensee et al. highlight that not all datasets are equally suitable for model comparison due to a high statistical variance (intra-method) or a low systematic (inter-method) variance. Suitable datasets, such as AMOS22, KiTS23, and LiTS, showed higher inter-method than intra-method standard deviation, whereas worse-suited datasets like BTCV showed lower standard deviation in their experiments. Improved convergence via PGPS-Performance yields reduced stochastic difference due to sampling, leading to improved segmentation performance and lower segmentation performance variance as seen in experiment 4.6. This provides a better signal-to-noise ratio, making PGPS-Performance a more reliable strategy for fair model comparison than standard CPS. Following the strategy in [32], we exemplarily compute inter- and intra-method standard deviation for PGPS-Performance for BTCV and KiTS23. Details on the experiment and results are reported in D. We found that PGPS-Performance sampling yields better method comparison, as the inter-method to intra-method standard deviation ratio is improved for both datasets when switching from CPS to PGPS-Performance sampling.
7 Conclusion and Outlook
In this work, we introduced a novel curriculum for semantic segmentation based on increasing the patch size during training. We evaluate the curriculum on 15 diverse and popular 3D medical image segmentation datasets against the default constant patch size baseline. We propose two curriculum modes: PGPS-Performance, which consistently outperforms constant patch size sampling on all 15 datasets, in terms of Dice Score, while simultaneously requiring only 90% of the original training time, and PGPS-Efficiency, which matches the performance of constant patch size training while drastically reducing training time to 44%. We demonstrate that modifying the sampling strategy substantially accelerates training convergence and statistically significantly improves final Dice score performance in deep learning for patch-based 3D medical image segmentation.
Progressive Resolution, a well-known input-length automatic curriculum based on image resolution [20, 21, 23], does lead to performance decreases in our experiments and is more suited to solve global image-level tasks like classification or contrastive pretraining. This underlines the need and research gap for a curriculum learning approach for dense prediction tasks. In contrast, our proposed curriculum based on patch size did improve the segmentation performance and is, to the best of our knowledge, the first input-length curriculum based on patch size in the computer vision and medical imaging domains.
Our analysis shows that tasks with strong class imbalance benefit most from PGPS, underscoring the importance of sampling strategies in patch-based training. Importantly, the advantages of PGPS are not limited to UNet-style architectures. We demonstrated consistent improvements for PGPS-Performance across both convolutional and transformer-based models, including UNet, UNETR, and SwinUNETR, highlighting the broad applicability of the method.
A key strength of our approach is its simple integration into any segmentation backbone training. Depending on the application, users can choose between PGPS-Efficiency for rapid experimentation with reduced compute and carbon footprint, or PGPS-Performance for maximizing segmentation performance in deployment settings, while also benefiting from a reduced carbon footprint. Furthermore, PGPS-Performance enables more reliable model comparison by improving the signal-to-noise ratio, by improving convergence, and thereby providing a clearer signal for methodological benchmarking.
In summary, we recommend PGPS-Performance as the new default sampling strategy for training 3D patch-based segmentation backbones for deployment. It is conceptually simple, easy to implement within existing frameworks such as nnU-Net, and provides consistent improvements in both segmentation performance and resource efficiency, establishing a strong novel sampling strategy in patch-based 3D medical image segmentation.
We see several promising directions for further research. The curriculum approach sequence length warmup was shown to reduce the gradient variance in training large language models, enabling larger learning rates and batch sizes, resulting in extremely short training time to only 6% of the original time [16]. We did not explore hyperparameter tuning in our experiments, but this may potentially also result in further runtime reduction.
While most research in medical image segmentation is predominantly conducted in the architectural design of backbones, we argue that the sampling strategy itself is a vastly overlooked area and requires more exploration. Advances in differentiable and adaptive sampling strategies, such as differentiable top-$k$ Patch Sampling [42], open new directions for optimizing patch-based sampling. Beyond patch sampling, alternative paradigms such as multi-scale frameworks [30] or cascaded architectures [2] aim to reduce reliance on patching altogether, but patch-based methods remain dominant in segmentation performance today [42].
Acknowledgements
SMF, JK and JAS acknowledge funding from Munich Center for Machine Learning (MCML). SMF, JCP, and JAS acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 515279324 / SPP 2177. SMF acknowledges funding from the EVUK program (”Next-generation AI for Integrated Diagnostics”) of the Free State of Bavaria. JK and JCP acknowledge funding from the Wilhelm Sander Foundation in Cancer Research (2022.032.1). JK acknowledges funding from the DAAD program Konrad Zuse Schools of Excellence in Reliable Artificial Intelligence, sponsored by the German Federal Ministry of Education and Research. RO and KL acknowledge funding from the EU Horizon Europe and Horizon 2020 research and innovation programme under grant agreement No 101057699 (RadioVal) and No 952103 (EuCanImage), respectively. RO acknowledges a research stay grant from the Helmholtz Information and Data Science Academy (HIDA). JAS and LD are supported by the German Federal Ministry of Research, Technology and Space (DECIPHER-M, 01KD2420G). JAS and DML acknowledge funding from HELMHOLTZ IMAGING, a platform of the Helmholtz Information & Data Science Incubator.
Data Availability
All data are publicly available. We provide detailed descriptions and proper citations for each dataset used to ensure full transparency and reproducibility. Additionally, all code for data preprocessing and preparation is publicly released and can be used to fully reproduce our results. Our code is publicly available at https://github.com/StefanFischer/nnUNet-ProgressiveGrowingOfPatchSize/releases/tag/v2.6.2.
Conflicts of Interest
The authors declare that they have no known conflict of interest nor competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the author(s) used ChatGPT in order to improve the readability of the manuscript. After using this tool, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
CRediT authorship contribution statement
Stefan M. Fischer: Conceptualization, Methodology, Software, Writing - Original Draft Johannes Kiechle: Validation, Writing - Review & Editing Laura Daza: Validation, Writing - Review & Editing Lina Felsner: Validation, Writing - Review & Editing, Richard Osuala: Validation, Writing - Review & Editing, Daniel Lang: Validation, Writing - Review & Editing, Karim Lekadir: Validation, Writing - Review & Editing, Funding acquisition, Jan C. Peeken: Validation, Writing - Review & Editing, Supervision, Funding acquisition, Julia A. Schnabel: Validation, Writing - Review & Editing, Supervision, Funding acquisition
Appendix A Batch Construction Strategies
Table 4: Comparison of Dice scores, FLOPs, and relative runtime across three datasets for Constant Patch Size (CPS) and Progressive Growing of Patch Size (PGPS) variants with split- and single-crop sampling. All values are normalized to the CPS baseline. PGPS-Perf (Single-Crop) also improves Dice score over CPS but incurs substantial runtime overhead due to increased dataloading costs.
Metric MSD6 Lung Tumor KiTS23 AMOS22
CPS PGPS-Perf (Split-Crop) PGPS-Perf (Single-Crop) CPS PGPS-Perf (Split-Crop) PGPS-Perf (Single-Crop) CPS PGPS-Perf (Split-Crop) PGPS-Perf (Single-Crop)
Dice Score [%] $\uparrow$ 70.00 72.77 72.45 86.02 86.46 87.18 88.62 88.78 88.77
Rel. FLOPs [%] $\downarrow$ 100 89 89 100 84 84 100 89 89
Rel. Runtime [%] $\downarrow$ 100 82 87 100 86 205 100 87 198
Setup: To efficiently utilize GPU resources during early PGPS-Performance stages with smaller patches, the batch size was increased either by (1) drawing patches from more patient volumes or (2) extracting multiple patches per patient volume. We investigated four patch sampling strategies, Single-Volume, Multi-Crop, Split-Crop, and Single-Crop, on the MSD Lung Tumor dataset to analyze their influence on convergence and generalization.
Batching Strategies: Single-Volume crops all patches from one patient volume, while Multi-Crop and Split-Crop sample from two patient volumes. In Split-Crop, one patient provides foreground patches and the other background patches, whereas Multi-Crop draws both from each patient. Single-Crop follows the nnU-Net scheme, loading as many patients as the batch size, drawing one patch per patient, thereby maximizing patient diversity, preventing patch overlaps.
Figure 6: Convergence of different batching strategies for PGPS-Performance on the MSD Lung Tumor dataset. Increasing batch size can be achieved by sampling from more patients or generating multiple crops per patient. Overfitting occurs when both foreground and background patches are drawn from the same patient.
Results on MSD Lung: All PGPS-Performance batching strategies converged faster than the baseline Constant Patch Size (CPS) setup (Fig. 6). Single-Volume exhibited strong overfitting, and Multi-Crop showed moderate overfitting. Split-Crop achieved the most stable convergence and highest Dice performance, while Single-Crop slightly lagged in convergence speed. Runtime analysis showed that Split-CropSplit-Crop and Multi-Crop ran efficiently (82% of CPS runtime), whereas Single-Crop required more dataloading time (87% of CPS runtime) because for each batch, many distinct patient volumes have to be loaded.
Cross-Dataset Runtime Comparison: While Split-Crop allows patch overlaps, Single-Crop prevents this. To quantify the effect on large high-quality benchmark datasets, we pick AMOS22 and KiTS23 to evaluate segmentation performance and computational costs. Table 4 reports Dice scores, FLOPs, and runtime normalized to the CPS baseline. Both PGPS variants consistently improved segmentation performance over CPS, but Split-Crop was much more efficient. Single-Crop runtime increased drastically, 205% on KiTS23 and 198% on AMOS22, caused by extensive I/O from loading large numbers of patient volumes. In contrast, Split-Crop achieved nearly identical performance with far lower computational overhead (82–87% of CPS runtime), while achieving higher segmentation performance than Single-Crop in AMOS22 and MSD Lung. For larger-batch datasets such as TotalSegmentatorV2 (batch size 1575), these runtime penalties for Single-Crop would become prohibitively large.
Conclusion: The choice of batch construction has a strong impact on generalization and computational efficiency. Split-Crop offers the best trade-off between convergence speed, segmentation performance, and runtime, and was therefore adopted as the default batching strategy for all PGPS-Performance experiments. The Single-Crop strategy remains useful for applications prioritizing maximal patch variety, though it would benefit from optimized dataloading implementations. Enlarging patch diversity by preventing patch overlaps did not seem to improve performance in our experiments.
Furthermore, we hypothesize that the low batch diversity is the main reason for overfitting. With batch diversity, we refer to the number of possible combinations between foreground and background patches of different patients. Given a dataset with $n$ samples, for Single-Volume only $n$ combinations are possible, as we pair foreground and background of Patient A. This low batch diversity ends up in strong overfitting.
For Multi-Crop, we pair the foreground and background of Patient A with those of Patient B. We thus build unordered pairs. For this strategy we only observe slight overfitting.
For Split-Crop, we pair the foreground of Patient A with the background of Patient B, which are ordered pairs and thus more as for Multi-Crop. This batch diversity is the same as for standard nnU-Net CPS training (batch size of two). Here, we do not observe any overfitting.
Maximal variety is possible with Single-Crop. As for each patch, a new patient is used, and thus, the maximal possible batch diversity. This large batch diversity did not improve performance over Split-Crop as there was slower convergence. Longer training might yield better results for Single-Crop than for Split-Crop.
Appendix B Orientation Sensitivity
Setup: Smaller patches reduce global spatial context, which can hinder segmentation of directionally defined structures (e.g., left/right or upper/lower). Orientation sensitivity varies across datasets: BTCV and AMOS22 each include 4 directionally defined classes out of 14 and 15, TotalSegmentatorV2 includes 66 out of 107, and ToothFairy2 includes 32 out of 42. Mirror augmentation, the nnU-Net default, can further obscure orientation cues by inverting anatomical structures. We trained CPS and PGPS-Performance with and without mirror augmentation on these datasets using 5-fold cross-validation.
Results: As shown in Table 5, the influence of mirror augmentation depends strongly on dataset orientation sensitivity. For BTCV and AMOS22, effects were minimal, with PGPS-Performance consistently outperforming CPS. In TotalSegmentatorV2, disabling mirroring substantially improved results, and PGPS-Performance achieved the best overall Dice score.
The strongest degradation occurred in ToothFairy2, where most classes are directionally defined; mirror augmentation severely reduced performance for CPS and PGPS-Performance, but particularly for the curriculum. Disabling mirroring did result in a large performance gain for both sampling strategies in ToothFairy2, where PGPS-Performance did outperform CPS.
Overall, PGPS-Performance remains beneficial when orientation-sensitive classes form only a subset of labels, but can suffer when such classes dominate the dataset. We therefore recommend caution with mirror augmentation in tasks with large numbers of directional-defined classes, as it may conflict with PGPS’s progressive curriculum and limit performance gains. Additionally, a CPS-trained network would also benefit from disabling mirroring.
Table 5: Impact of mirror augmentation on CPS and PGPS across datasets containing directionally defined semantic classes. Results are mean Dice Scores (%) over 5-fold cross-validation.
Dataset With Mirror Without Mirror
CPS PGPS-Perf CPS PGPS-Perf
BTCV 83.37 83.81 82.93 83.09
AMOS22 88.62 88.78 88.56 88.63
TotalSegV2 84.56 84.64 87.82 88.15
ToothFairy2 63.23 47.08 76.92 77.00
Appendix C nnU-Net Performance in Related Work
Table 6: Comparison of nnU-Net baselines and related models across the Medical Segmentation Decathlon (MSD), BTCV, KiTS23, and AMOS22-CT/MRI datasets. Reported values denote mean Dice Scores over 5-fold cross-validation as provided in the respective publications. All models, except the last row, were trained using the default constant patch size (CPS) strategy. The final row reports performance with our proposed Progressive Growing of Patch Size (PGPS-Performance) curriculum. Bold values indicate the best result per dataset. This overview is likely not exhaustive. Overall, Models trained via our proposed PGPS-Performance curriculum yield the best segmentation performance, delivering the highest performance on all datasets, except MSD2.
Original Publication MSD1 MSD2 MSD3 MSD4 MSD5 MSD6 MSD7 MSD8 MSD9 MSD10 BTCV KiTS23 AMOS22-CT/MRI
Constant Patch Size:
nnU-Net (orig.) [29] 74.11 93.28 79.71 88.91 75.37 72.11 67.45 68.37 96.38 45.53 82.79
nnU-Net v2 (Repo) [43] 73.98 93.34 79.53 88.95 75.01 69.83 67.17 68.41 96.90 45.04 83.25
nnU-Net revisited [32] 83.08 86.04 88.64
MedNeXt [27] 83.56
RecycleNet [44] 82.96 88.58
Extending nnU-Net is All You Need [45] 88.64
Auto-nnU-Net [46] 73.98 93.39 79.45 89.04 73.53 68.33 66.07 68.31 96.66 46.04
Ours (CPS) 74.12 93.29 78.74 88.96 73.13 70.00 68.68 68.61 97.02 48.41 83.37 86.02 88.62
Progressive Growing of Patch Size:
Ours (PGPS-Perf) 74.15 93.29 80.60 89.15 76.31 72.77 68.80 68.98 97.15 51.02 83.81 86.46 88.78
Setup: We assessed the variability in segmentation performance that was reported for nnU-Net baselines in the literature. Therefore, we reviewed publications that employed nnU-Net in their studies and reported results from 5-fold cross-validation on datasets contained in our benchmark.
Results: In total, we identified nine sources (eight publications and the official DKFZ nnU-Netv2 GitHub repository) that provided relevant validation scores. Although this collection is not exhaustive, it reflects the diversity of nnU-Net results currently cited in the field.
Table 6 summarizes Dice Scores reported across the Medical Segmentation Decathlon (MSD), BTCV, KiTS23, and AMOS22 datasets. Most of these results originate from the DKFZ research group [27, 29, 32, 43, 44, 45], who developed and maintain nnU-Net. Another external publication was found using nnU-Net as baseline [46]. In Figure 7, we plot the normalized standard deviation in performance over all found runs per dataset. Note that scores originate from different nnU-Net versions (nnU-Netv1 and nnU-Netv2), but DKFZ claims that both maintain equivalent segmentation performance [43].
For datasets, where we found multiple CPS-based results, as seen in Table 6, the variance in reported performance underscores the variability of training outcomes across different runs. Despite this variability, our proposed PGPS-Performance curriculum consistently achieves the highest Dice Score across all evaluated datasets except MSD2.
The marginal exception in MSD2 could result from convergence saturation, where other factors, such as network initialization or cached class sampling origins, might play a stronger role than improved sampling.
Figure 7: Reported nnU-Net Segmentation Performances in Literature compared to PGPS-Performance Sampling. We collected reported nnU-Net performances, trained via constant patch size (CPS), and plotted their normalized standard deviation in Dice score against instances trained via PGPS-Performance. Training with PGPS-Performance resulted in the best-performing models in all datasets, except for MSD2. For all datasets, we had 4 or more CPS performance values, while for KiTS23, we only had 2 values.
Overall, these findings confirm that our proposed PGPS-Performance does result in improved performance over repeated runs on different hardware and environments.
Appendix D More Reliable Backbone Comparison
Setup: As discussed in Section 6.4, the proposed PGPS-Performance strategy is expected to provide a more reliable basis for model comparison than standard Constant Patch Size (CPS) training. To evaluate this assumption, we conducted experiments on the BTCV and KiTS23 datasets, which were identified by [32] as the least and most reliable benchmarks for backbone comparison, respectively. Following [32], we compute the ratio between the inter-method and intra-method standard deviations. A higher ratio indicates that architectural differences dominate over random variation, meaning that the dataset and training setup allow for more meaningful model comparison. We used the nnU-Net ResEnc variants (M, L, and XL), which scale both patch size and model capacity, alongside the original nnU-Net configuration. Each model was trained using both CPS and PGPS-Performance sampling strategies, allowing us to assess whether PGPS-Performance improves the signal-to-noise ratio in backbone comparisons.
Results: Table 7 summarizes mean Dice scores across both datasets. For BTCV, the lowest performing nnU-Net variant trained via PGPS-Performance resulted in even higher segmentation performance than all model variants for CPS sampling. Also for KiTS23, PGPS-Performance sampling improved segmentation performance over CPS sampling, while increasing model capacity had a higher impact on the segmentation performance for all variants.
Table 7: Comparison of Dice Scores between Constant Patch Size (CPS) and Progressive Growing of Patch Size (PGPS) training strategies across nnU-Net and ResEnc variants on BTCV and KiTS23 datasets. Values indicate mean Dice Scores over 5-fold cross-validation. Bold values mark the highest performance per architecture and dataset.
Model BTCV KiTS23
CPS PGPS-Perf CPS PGPS-Perf
nnU-Net (orig.) 83.37 83.81 86.02 86.46
ResEncM 83.48 83.97 87.05 88.00
ResEncL 83.30 84.18 88.24 88.64
ResEncXL 83.30 84.03 88.68 88.93
While PGPS-Performance generally improved absolute performance for all architectures, another critical finding concerns the reduction in statistical variance across training runs and architectures. Table 8 shows the intra- and inter-method standard deviations and the resulting inter/intra ratios. On the KiTS23 dataset, previously shown to provide the very stable comparison signal [32], the ratio remains nearly constant (76.4% vs. 76.7%), indicating that both CPS and PGPS-Performance allow for consistent differentiation of model scales. In contrast, on BTCV, which is known for a low signal-to-noise ratio for backbone comparison, PGPS-Performance substantially increased the ratio from 32.6% to 55.6%. This demonstrates that PGPS-Performance effectively increases the signal-to-noise ratio, thereby yielding more reliable and interpretable backbone comparisons. Intra-Method SD is only decreased for KiTS23, while for BTCV it is increased. In our experiment in section 4.6, we focused on repeating training on the same fold. In contrast, here the standard deviation is computed over five folds, such that fold difficulty (annotation errors, sample difficulty) itself also plays an important role.
Overall, these findings confirm that PGPS-Performance not only enhances segmentation performance but also improves the experimental robustness of comparative studies. By improving training convergence, the proposed curriculum increases the clarity of architectural performance differences.
Table 8: Comparison of Dice score variability across nnU-Net and ResEnc variants for Constant Patch Size (CPS) and Progressive Growing of Patch Size (PGPS) on BTCV and KiTS23. Reported values represent the mean standard deviation of the Dice score over 5-fold cross-validation. The inter/intra-method ratio indicates how clearly the dataset ranks architectures (higher is better). PGPS-Performance increases this ratio, improving backbone comparability.
BTCV KiTS23
CPS PGPS-Perf CPS PGPS-Perf
nnU-Net (orig.) 2.24 2.54 1.77 1.73
ResEncM 2.09 2.25 1.55 1.40
ResEncL 2.52 2.48 0.92 0.94
ResEncXL 2.20 2.27 1.19 0.91
Avg. Intra-Method SD [%] 2.27 2.38 1.358 1.245
Inter-Method SD [%] 0.74 1.33 1.037 0.955
Inter/Intra Ratio [%] $\uparrow$ 32.6 55.6 76.4 76.7